In a study, 31% of adults questioned. Among 13 randomly selected, only 1 reported that their health was excellent.
a) Find the Probability that when 13 adults are randomly selected, exactly 1 is in excellent health.
b) Find the probability that when 13 adults are randomly selected, at most 1 is in excellent health.
I'm having trouble on trying to figure out what formula to use for these questions.. and how to remember which ones to use for questions like these.
Thank you.

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

binomial function n=13 x is number reported and p=0.31 with 1-p=0.69
for 1 exactly it is 13C1*0.31^1*0.69^12, the 13C1=13 ways it happens
this is 0.0197
for at most 1 is the above + probability of 0's occurring or .69^13=0.0080
that probability is 0.0277
Notice how the formula is structured and how the exponents add up to n.

Answer by ikleyn(52754)

.
In a study, 31% of adults questioned. Among 13 randomly selected, only 1 reported that their health was excellent.

a) Find the Probability that when 13 adults are randomly selected, exactly 1 is in excellent health.

b) Find the probability that when 13 adults are randomly selected, at most 1 is in excellent health.

I'm having trouble on trying to figure out what formula to use for these questions.. and how to remember which ones to use for questions like these.

Thank you.
~~~~~~~~~~~~~~~


Actually,  as this problem is worded/formulated in the post,  it is  HEAVILY  SICK.


In order for the formulation be correct,  it should be formulated  DIFFERENTLY.


If to treat it literally as it is written,  it is partly self-contradictory and partly nonsensical.


The sum of these parts is equal to the whole thing,  i.e.  100%.


The diagnosis is that in this form,  the problem  CAN  NOT  EXIST.


In this form,  it is not a  Math  problem.   It is a  FAKE,  instead.